# LiNo: Advancing Recursive Residual Decomposition of Linear and Nonlinear Patterns for Robust Time Series Forecasting

### v1

@@ 4 Experiments
To thoroughly evaluate the performance of our proposed LiNo, we conduct extensive experiments on 13 widely used, real-world datasets including ETT (4 subsets) (Zhou et al., 2022a), Traffic, Exchange, Electricity(ECL), Weather (Wu et al., 2021), Solar-Energy(Solar) (Lai et al., 2018) and PEMS (4 subsets) (LIU et al., 2022). Detailed descriptions of the datasets can be found in Appendix A. We select both univariate and multivariate time series forecasting tasks, ensuring a comprehensive assessment.
@@
All the experiments are conducted on a single NVIDIA GeForce RTX 4090 with 24G VRAM. The mean squared error (MSE) loss function is utilized for model optimization. We use the ADAM optimizer with an early stop parameter $patience=6$. To foster reproducibility, we make our code, training scripts, and some visualization examples available in this GitHub Repository [^1] https://github.com/Levi-Ackman/LiNo. Full implementation details and other information can be found in Appendix B.1.
@@ 4.1 Multivariate Time Series Forecasting Results
Table 1: Multivariate forecasting results with prediction lengths $F\in\{12,24,36,48\}$ for PEMS dataset while $F\in\{96,192,336,720\}$ for others with fixed lookback window $T=96$. Results are averaged from all prediction lengths. Avg means further averaged by subsets. Full result is left in Appendix E.1 due to space limit.
@@ 4.2 Univariate Time Series Forecasting Results
Table 2: Univariate forecasting results with prediction lengths $F\in\{96,192,336,720\}$ and fixed lookback length $T=96$ for all datasets. Results are averaged from all prediction lengths. Full result is left in Appendix E.2 due to space limit.
@@ 4.3 LiNo Analysis
Increasing lookback length.
It is generally expected that increasing the input length will enhance forecasting performance by incorporating more information (Zeng et al., 2023). This improvement is typically observed in linear forecasts, supported by statistical methods (Liu et al., 2023) that utilize extended historical data. Figure 3 evaluates the performance of LiNo and other prestigious baselines. LiNo effectively leverages longer lookback windows, showing a positive correlation between predictive performance and input length. It significantly outperforms other baselines on the Weather and ECL datasets and achieves comparable results on the ETTm2 dataset. Further analysis of LiNo is provided in Appendix C.
@@ 4.4 Analysis of Different Forecasting Model Designs
We use iTransformer, a leading transformer-based time series forecasting model, as the backbone. We evaluate three model designs: ‘Raw’ (classical design), ‘Mu’ (recursive splitting of representations for prediction), and ‘LiNo’ (our proposed framework with further recursive splitting into linear and nonlinear patterns). The forecasting performance in Table 5 (b) demonstrates the effectiveness of the LiNo framework. Compared to ‘Raw’, ‘Mu’ reduces the MSE on ETTm2, ECL, and Weather by 1.3%, 0.28%, and 1.06%, respectively. LiNo further reduces the MSE by 2.96%, 6.34%, and 6.72%. These results indicate that while ‘Mu’ improves forecasting performance, it remains suboptimal due to the entanglement of linear and nonlinear predictions. LiNo effectively separates these patterns, achieving more accurate results.
@@
Visualization of linear and nonlinear predictions.
Figure 5: Visualization of multivariate forecasting result (last channel/variate) of the proposed LiNo on ETTh1 and ECL datasets. We set the number of LiNo blocks (layers) to $N=3$. The task is multivariate time series forecasting with input $T=96$ and target $F=96$. ‘LP’ denotes Linear prediction, and ‘NP’ stands for Nonlinear prediction. LP $i$ or NP $i$ ($i\in\{1,2,3\}$) is the linear or nonlinear prediction of $i$-th layer (level).
We visualize LiNo’s forecasting results on the ETTh1 and ECL datasets in Figure 5. The Li block primarily captures linear patterns such as long-term trends, while the No block effectively captures nonlinear signals like fluctuations and seasonality. The forecasting results for ECL and ETTh1 reveal three distinct linear modes and three different nonlinear patterns, enhancing the interpretability of time series forecasting and aiding in understanding the underlying data dynamics. Additional visualization results of LiNo can be found in Appendix D.
@@ 5 Conclusion
6 Ethics Statement
Our work only focuses on the scientific problem, all datasets are publicly available, so there is no potential ethical risk.
7 Reproducibility Statement
We involve the implementation details in Appendix B. The source code is accessible in GitHub (https://github.com/Levi-Ackman/LiNo) for reproducibility.
@@ Appendix A Datasets Description
    ETT (Electricity Transformer Temperature) Zhou et al. (2022a)  [^2] https://github.com/zhouhaoyi/ETDataset comprises two hourly-level datasets (ETTh) and two 15-minute-level datasets (ETTm). Each dataset contains seven oil and load features of electricity transformers from July $2016$ to July $2018$.
@@
    Solar-Energy Lai et al. (2018)  [^7] https://github.com/thuml/iTransformer records the solar power production of 137 PV plants in 2006, which is sampled every 10 minutes.
@@
    PEMS (LIU et al., 2022)  [^8] https://pems.dot.ca.gov/ contains public traffic network data in California collected by 5-minute windows.
@@ B.1 Experiment Details
To foster reproducibility, we make our code, training scripts, and some visualization examples available in this GitHub Repository [^9] https://github.com/Levi-Ackman/LiNo. All the experiments are conducted on a single NVIDIA GeForce RTX 4090 with 24G VRAM. The mean squared error (MSE) loss function is utilized for model optimization. We use the ADAM optimizer with an early stop parameter $patience=6$. We explore the number of LiNo blocks $N\in\{1,2,3,4\}$, dropout ratio $dp\in\{0.0,0.2,0.5\}$, and the dimension of layers $dim\in\{256,512\}$. The $learning\ rate\in\{1e-3,1e-4,1e-5\}$ and $batch\ size\in\{32,64,128,256\}$ are adjusted based on the size and dimensionality of the dataset, as well as the specific conditions of our experimental setup. All the compared multivariate forecasting baseline models that we reproduced are implemented based on the benchmark of Time series Lab (Wang et al., 2024b) Repository, which is fairly built on the configurations provided by each model’s original paper or official code.
@@ C.2 Model Efficiency
Appendix D Visualization
D.1 Visualization of LiNo’s Forecasting Results

### v4

@@ 4 Experiments
To thoroughly evaluate the performance of our proposed LiNo, we conduct extensive experiments on 13 widely used, real-world datasets including ETT (4 subsets) (Zhou et al., 2022a), Traffic, Exchange, Electricity(ECL), Weather (Wu et al., 2021), Solar-Energy (Solar) (Lai et al., 2018) and PEMS (4 subsets) (Liu et al., 2022a). Detailed descriptions of the datasets can be found in Appendix A. We select univariate and multivariate time series forecasting tasks, ensuring a comprehensive assessment.
@@
All the experiments are conducted on a single NVIDIA GeForce RTX 4090 with 24G VRAM. The mean squared error (MSE) loss function is utilized for model optimization. We use the ADAM optimizer with an early stop parameter $patience=6$. To foster reproducibility, our code, training scripts, and some visualization examples are available in this GitHub Repository [^1] https://github.com/Levi-Ackman/LiNo. Full implementation details and other information are in Appendix B.1.
@@ 4.1 Multivariate Time Series Forecasting Results
Table 1: Multivariate forecasting results with prediction lengths $F\in\{12,24,36,48\}$ for PEMS dataset while $F\in\{96,192,336,720\}$ for others with fixed lookback window $T=96$. Results are averaged from all prediction lengths. Avg means further averaged by subsets. Full result is left in Appendix H.1 due to space limit.
@@ 4.2 Univariate Time Series Forecasting Results
Table 2: Univariate forecasting results with prediction lengths $F\in\{96,192,336,720\}$ and fixed lookback length $T=96$ for all datasets. Results are averaged from all prediction lengths. Full result is left in Appendix H.2 due to space limit.
@@ 4.3 LiNo Analysis
Further ablation on No Block.
Table 5: Further ablation study on the No Block. ‘Temporal’ means only extracting temporal variations (without introducing nonlinearity), ‘Frequency’ strands for frequency information extraction (without introducing nonlinearity), and ‘TF’ means these two patterns are captured simultaneously. The nonlinearity of the No Block is introduced by adding ReLU or Tanh activation function.
Models Temporal Frequency TF ReLU Tanh
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.385 0.398 0.383 0.393 0.383 0.392 0.384 0.398 0.375 0.394
720 0.487 0.475 0.482 0.469 0.478 0.467 0.488 0.477 0.464 0.458
ETTm1 96 0.341 0.369 0.342 0.369 0.338 0.366 0.328 0.365 0.322 0.359
720 0.481 0.447 0.479 0.445 0.474 0.439 0.478 0.445 0.465 0.442
ECL 96 0.186 0.268 0.187 0.267 0.183 0.263 0.154 0.248 0.150 0.243
720 0.245 0.320 0.245 0.320 0.241 0.317 0.223 0.311 0.221 0.308
Considering that in real-world scenarios, nonlinear temporal variations and frequency information are more prevalent than inter-series dependencies in most cases, we conducted further ablation studies. Specifically, without considering inter-series dependencies, we simplified the No block to retain only the Temporal&Frequency Extraction component (TF_Extraction) shown in Figure 2 (left), and performed additional ablation experiments.
Combining Temporal and Frequency projections enhances expressiveness, improving the model’s ability to learn complex patterns. Tanh provides a smooth, symmetric activation function that better captures the complex dependencies in both time and frequency domains, therefore delivering a better performance than the more commonly used ReLU, as in Table 5.
@@
Further analysis of LiNo is provided in Appendix C to save space.
@@ 4.4 Analysis of Different Forecasting Model Designs
We use iTransformer, a leading transformer-based time series forecasting model, as the backbone. We evaluate three model designs: ‘Raw’ (classical design), ‘Mu’ (recursive splitting of representations for prediction, design used in N-BEATS (Oreshkin et al., 2020)), and ‘LiNo’ (our proposed framework with further recursive splitting into linear and nonlinear patterns). The forecasting performance in Table 6 (b) demonstrates the effectiveness of the LiNo framework. Compared to ‘Raw’, ‘Mu’ reduces the MSE on ETTm2, ECL, and Weather by 1.3%, 0.28%, and 1.06%, respectively. LiNo further reduces the MSE by 2.96%, 6.34%, and 6.72%. These results indicate that while ‘Mu’ improves forecasting performance, it remains suboptimal due to the entanglement of linear and nonlinear predictions. LiNo effectively separates these patterns, achieving more accurate results.
@@
We leave some important visualization results of LiNo, such as model Weight and Prediction of Li&No Block, in Appendix G.
@@ 5 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Datasets Description
    ETT (Electricity Transformer Temperature) (Zhou et al., 2022a)  [^2] https://github.com/zhouhaoyi/ETDataset comprises two hourly-level datasets (ETTh) and two 15-minute-level datasets (ETTm). Each dataset contains seven oil and load features of electricity transformers from July $2016$ to July $2018$.
@@
    Solar-Energy (Lai et al., 2018)  [^7] https://github.com/thuml/iTransformer records the solar power production of 137 PV plants in 2006, which is sampled every 10 minutes.
@@
    PEMS (Liu et al., 2022a)  [^8] https://pems.dot.ca.gov/ contains public traffic network data in California collected by 5-minute windows.
@@ B.1 Experiment Details
To foster reproducibility, our code, training scripts, and some visualization examples are available in this GitHub Repository [^9] https://github.com/Levi-Ackman/LiNo. All the experiments are conducted on a single NVIDIA GeForce RTX 4090 with 24G VRAM. The mean squared error (MSE) loss function is utilized for model optimization. We use the ADAM optimizer with an early stop parameter $patience=6$. We explore the number of LiNo blocks $N\in\{1,2,3,4\}$, dropout ratio $dp\in\{0.0,0.2,0.5\}$, and the dimension of layers $dim\in\{256,512\}$. The $learning\ rate\in\{1e-3,1e-4,1e-5\}$ and $batch\ size\in\{32,64,128,256\}$ are adjusted based on the size and dimensionality of the dataset, as well as the specific conditions of our experimental setup. All the compared multivariate forecasting baseline models that we reproduced are implemented based on the benchmark of Time series Lab (Wang et al., 2024b) Repository  [^10] https://github.com/thuml/Time-Series-Library, which is fairly built on the configurations provided by each model’s original paper or official code.
@@ C.2 Model Efficiency
C.3 Increasing lookback length
It is generally expected that increasing the input length will enhance forecasting performance by incorporating more information (Zeng et al., 2023). This improvement is typically observed in linear forecasts, supported by statistical methods (Liu et al., 2023) that utilize extended historical data. Figure 4 evaluates the performance of LiNo and other prestigious baselines. LiNo effectively leverages longer lookback windows, showing a positive correlation between predictive performance and input length. It significantly outperforms other baselines on the Weather and ECL datasets and achieves comparable results on the ETTm2 dataset.
Appendix D Comparison between LiNo and N-BEATS
D.1 Theory analysis of N-BEATS and LiNo
Table 10: Ablations on overall design under Univariate Scenario with input length fixed to $T=96$. ‘RAW’ refers to the raw model. ‘LN’ stands for separately modeling linear and nonlinear patterns, but no decomposition is applied. ‘N-BEATS’ refers to the N-BEATS-style model where RRD is employed but without separating the extraction and prediction of linear and nonlinear. ‘LiNo’ is the proposed framework, where both RRD and separation of linear and nonlinear patterns. The bold values indicate the best performance.
Models RAW LN N-BEATS LiNo
Metric MSE MAE MSE MAE MSE MAE MSE MAE
ETTh2 96 0.133 0.282 0.128 0.274 0.126 0.271 0.126 0.271
192 0.183 0.335 0.177 0.327 0.179 0.328 0.174 0.324
336 0.216 0.372 0.213 0.367 0.213 0.368 0.209 0.364
720 0.227 0.384 0.229 0.385 0.226 0.381 0.223 0.379
ETTm2 96 0.067 0.186 0.068 0.188 0.067 0.187 0.066 0.186
192 0.102 0.239 0.102 0.237 0.102 0.238 0.099 0.234
336 0.132 0.276 0.133 0.274 0.132 0.276 0.130 0.272
720 0.185 0.334 0.185 0.334 0.185 0.333 0.187 0.335
Traffic 96 0.167 0.247 0.163 0.255 0.168 0.253 0.159 0.248
192 0.164 0.245 0.158 0.244 0.159 0.244 0.155 0.241
336 0.163 0.247 0.158 0.245 0.156 0.242 0.153 0.237
720 0.183 0.263 0.176 0.261 0.177 0.262 0.174 0.260
Avg 0.160 0.284 0.158 0.283 0.158 0.282 0.155 0.279
To better assist the reader in understanding the similarities and differences between our work and N-BEATS (Oreshkin et al., 2020), and to help clarify our contributions, we have provided the following analysis.
Here are some essential definitions:
  1. 1.
    The linear feature extractor is denoted as $Li\_Block$, the extracted pattern $L_{i}$, and linear prediction $P^{li}_{i}$.
  2. 2.
    The nonlinear feature extractor is denoted as $No\_Block$, the extracted pattern $N_{i}$, and nonlinear prediction $P^{no}_{i}$.
  3. 3.
    The projection to get the prediction is denoted as $FC$
  4. 4.
    We assume the input to the current layer in deep neural network as $H_{i}$ and the next layer $H_{i+1}$, where $H_{0}=X\_embed=XW+b$, $X$ is the input raw time series
  5. 5.
    Prediction of current layer $P_{i}$
  6. 6.
    Final prediction $\hat{Y}$
Then, N-BEATS can be summarized as:
$N_{i}$ $=No\_Block(H_{i}),$ (7)
$P_{i}$ $=FC(N_{i}),$
$H_{i+1}$ $=H_{i}-N_{i},$
$\hat{Y}$ $=\sum_{i=1}^{N}P^{i}.$
Our LiNo can be formulated as:
$L_{i}$ $=Li\_Block(H_{i}),$ (8)
$P^{li}_{i}$ $=FC(L_{i}),$
$R^{L}_{i}$ $=H_{i}-L_{i},$
$N_{i}$ $=No\_Block(R^{L}_{i}),$
$P^{no}_{i}$ $=FC(N_{i}),$
$R^{N}_{i}$ $=R^{L}_{i}-N_{i},$
$H_{i+1}$ $=R^{N}_{i},$
$P^{i}$ $=P^{li}_{i}+P^{no}_{i},$
$\hat{Y}$ $=\sum_{i=1}^{N}P^{i}.$
Below are some differences:
  1. 1.
    LiNo introduces a two-stage extraction process: linear ($Li\_Block$) and nonlinear ($No\_Block$), while N-BEATS only uses a one-stage extraction and without separation of linear and nonlinear extraction.
  2. 2.
    LiNo computes predictions from both linear and nonlinear components, while N-BEATS only uses the nonlinear output.
  3. 3.
    LiNo refines intermediate representations after each extraction stage, whereas N-BEATS directly modifies hidden states.
  4. 4.
    LiNo combines both linear and nonlinear predictions, while N-BEATS only obtains nonlinear predictions.
D.2 Experiment results analysis of different model design
We designed the following experiments to validate the superiority of LiNo over N-BEATS-style design under a Univariate Scenario, to exclude any interference from channel-independent or channel-dependent information.
The learnable autoregressive model (AR) mentioned in the paper is employed as the linear feature extractor ($Li\_Block$).
The Temporal + Frequency projection, combined with the Tanh activation function, is adopted as the nonlinear feature extractor ($No\_Block$).
To ensure a fair comparison with the N-BEATS-style design (since completely removing the $Li\_Block$ may result in potential loss of parameters), the $No\_Block$ used in N-BEATS is actually a combination of $Li\_Block$ and $No\_Block$, which is still nonlinear:
$L_{i}$ $=Li\_Block(H_{i}),$ (9)
$N_{i}$ $=No\_Block(L_{i}),$
$P_{i}$ $=FC(N_{i}),$
$H_{i+1}$ $=H_{i}-N_{i},$
$\hat{Y}$ $=\sum_{i=1}^{N}P^{i}.$
We also include two more designs to investigate the specific effectiveness of LiNo.
RAW.
The input features pass through the Li Block and No Block sequentially, but no feature decomposition is performed. The model’s final output features are directly used for prediction (Traditional Design).
$L_{i}$ $=Li\_Block(H_{i}),$ (10)
$N_{i}$ $=No\_Block(L_{i}),$
$H_{i+1}$ $=N_{i},$
$\hat{Y}$ $=FC(H_{N}).$
LN.
The input features pass through the Li Block and No Block sequentially, with each block generating its own prediction based on the features extracted, but no residual decomposition is applied (Common Linear-Nonlinear Decomposition Design).
$L_{i}$ $=Li\_Block(H_{i}),$ (11)
$P^{li}_{i}$ $=FC(L_{i}),$
$N_{i}$ $=No\_Block(L_{i}),$
$P^{no}_{i}$ $=FC(N_{i}),$
$H_{i+1}$ $=N_{i},$
$\hat{Y}$ $=\sum_{i=1}^{N}(P^{li}_{i}+P^{no}_{i}).$
Experimental results in Table 10 demonstrate the superiority of LiNo over the N-BEATS-style design. Although both N-BEATS and LN show slight improvements over RAW, our LiNo design significantly outperforms the other designs. This demonstrates:
  1. 1.
    The effectiveness of explicit linear and nonlinear modeling.
  2. 2.
    The effectiveness of RRD for representation decomposition.
  3. 3.
    The potential for further improvement by combining both approaches (LiNo).
D.3 Comparison between LiNo and original N-BEATS
Table 11: Multivariate forecasting results with prediction lengths $F\in\{96,192,336,720\}$ and fixed lookback winodw $T=96$ for all datasets. The Results of N-BEATS and N-HiTS are taken from  (Challu et al., 2023). The bold values indicate the best performance.
Models
LiNo
N-HiTS
N-BEATS
Metric
MSE
MAE
MSE
MAE
MSE
MAE
ETTm2
96
0.171
0.254
0.176 0.255
0.184
0.263
192
0.237
0.298
0.245 0.305
0.273
0.337
336
0.296
0.336
0.295 0.346
0.309
0.355
720
0.395
0.393
0.401 0.413
0.411
0.425
Promote
-6.63%
-7.17%
-5.10% -4.42%
0
0
ECL
96
0.138
0.233
0.147
0.249
0.145
0.247
192
0.155
0.250
0.167 0.269
0.180
0.283
336
0.171
0.267
0.186 0.290
0.200
0.308
720
0.191
0.290
0.243 0.340
0.266
0.362
Promote
-17.19%
-13.33%
-6.07% -4.33%
0
0
Weather
96
0.154
0.199
0.158 0.195
0.167
0.203
192
0.205
0.248
0.211 0.247
0.229
0.261
336
0.262
0.290
0.274 0.300
0.287
0.304
720
0.343
0.342
0.351 0.353
0.368
0.359
Promote
-8.28%
-4.26%
-5.42% -2.84%
0
0
We also include the comparison of LiNo and the original version of N-BEATS. As in Table 11, LiNo significantly outperforms N-BEATS (Oreshkin et al., 2020) and its successor, N-HiTS (Challu et al., 2023), across the ETTm2, ECL, and Weather datasets. This demonstrates that, compared to using Recursive Residual Decomposition (RRD) to refine the final prediction based on residual errors from previous predictions (as proposed in N-BEATS), employing RRD to capture more nuanced linear and nonlinear patterns (as in this work) is a better choice.
Appendix E Comparison between proposed No block and SOFTS
Table 12: Multivariate forecasting performance comparison between SOFTS and the proposed No Block with input length fixed to $T=96$. The bold values indicate the best performance. The result of SOFTS is directly referenced from the original paper (Han et al., 2024).
Models ETTh1 ETTh2 ETTm1 ETTm2 Weather
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
SOFTS 96 0.381 0.399 0.297 0.347 0.325 0.361 0.180 0.261 0.166 0.208
192 0.435 0.431 0.373 0.394 0.375 0.389 0.246 0.306 0.217 0.253
336 0.480 0.452 0.410 0.426 0.405 0.412 0.319 0.352 0.282 0.300
720 0.499 0.488 0.411 0.433 0.466 0.447 0.405 0.401 0.356 0.351
No Block 96 0.377 0.394 0.293 0.342 0.333 0.365 0.175 0.257 0.172 0.211
192 0.423 0.420 0.372 0.390 0.369 0.385 0.241 0.300 0.214 0.249
336 0.457 0.440 0.414 0.425 0.407 0.407 0.302 0.340 0.276 0.296
720 0.454 0.457 0.417 0.438 0.470 0.440 0.399 0.396 0.353 0.347
We noticed that there is a recent work that is so close to our channel mixing technique, which is called SOFTS (NeurIPS 2024). Our channel mixing technique and SOFTS both stem from the idea of learning channel dependence while maintaining channel independence. Despite that, here are some major differences between LiNo and SOFTS:
  • LiNo uses the weighted sum of all channels with weight generated by the softmax function, while SOFTS uses a stochastic pooling technique to obtain the global token, which is more time-consuming.
  • We directly perform softmax to input feature, while SOFTS first go through a Feedforward Network, which is redundant.
Experiment results in Table 12 successfully demonstrate the superiority of the proposed No block over SOFTS.
Appendix F Compare with extensive baselines
Table 13: Multivariate forecasting results comparison between LiNo, TimeMixer, and SegRNN models with input length fixed to $T=96$. The bold values indicate the best performance for each dataset.
Models LiNo TimeMixer SegRNN
Metric MSE MAE MSE MAE MSE MAE
ETTm1 96 0.322 0.361 0.328 0.367 0.342 0.379
192 0.365 0.383 0.369 0.389 0.383 0.402
336 0.401 0.408 0.404 0.411 0.407 0.420
720 0.469 0.447 0.473 0.451 0.471 0.455
ETTm2 96 0.171 0.254 0.176 0.259 0.176 0.259
192 0.237 0.298 0.242 0.303 0.241 0.305
336 0.296 0.336 0.303 0.339 0.301 0.346
720 0.395 0.393 0.396 0.399 0.425 0.436
ETTh1 96 0.378 0.395 0.384 0.400 0.385 0.411
192 0.423 0.423 0.437 0.429 0.434 0.441
336 0.455 0.438 0.472 0.446 0.462 0.463
720 0.459 0.456 0.508 0.489 0.497 0.488
ETTh2 96 0.292 0.340 0.297 0.348 0.284 0.340
192 0.375 0.391 0.369 0.392 0.375 0.396
336 0.418 0.426 0.427 0.435 0.425 0.437
720 0.422 0.441 0.442 0.461 0.431 0.454
Weather 96 0.154 0.199 0.162 0.208 0.165 0.230
192 0.205 0.248 0.208 0.252 0.213 0.277
336 0.262 0.290 0.263 0.293 0.274 0.323
720 0.343 0.342 0.345 0.345 0.354 0.372
1st Count 19 20 0 0 1 0
After careful consideration, we additionally selected SegRNN (Lin et al., 2023) as an RNN-based baseline and considered another well-known work, TimeMixer (Wang et al., 2024a), which is also based on trend (linear) and seasonal (nonlinear) decomposition.
We still utilized the well-known Time series Lab (Wang et al., 2024b) Repository  [^11] https://github.com/thuml/Time-Series-Library.
LiNo still consistently outperforms overall, achieving leading results in 19/20 cases for MSE and 20/20 cases for MAE, as illustrated in Table 13.
Appendix G Visualization
G.1 Visualization of LiNo’s Forecasting Results